In 2012, the Campbell Soup Company acquired Bolthouse Farms for $1.55 billion. This acquisition increased the level of vertical integration in Campbell, as Bolthouse Farms owned and operated extensive farming operations where it produced many food items used in Campbell's products.  

Suppose that the value of produce provided by these farms after the acquisition would be

$90

million per year for Campbell and that, in addition, Campbell could save

$20

million per year in costs it would otherwise spend in searching for and negotiating over equivalent produce. However, the transaction cost of the acquisition (lawyers' fees, re-locating some production facilities, paying severance to unnecessary employees, etc.) required a one-time payment of

$50

million.  

Assume that Campbell receives the produce related benefits in perpetuity. If the interest rate used to discount future earnings is

5

%,

what is the present value of net benefits to Campbell?

The present value (PV) of net benefits to Campbell is
Answer #1

Initial cost ($ Million) = 1,550 + 50 = 1,600

Total annual savings ($ Million) = Annual value of produce + Annual savings = 90 + 20 = 110

PV of Total annual savings ($ Million) = 110 / 0.05** = 2,200

PV of net benefits ($ Million) = PV of Total annual savings - Initial cost = 2,200 - 1,600

= 600